# Memento: Towards Proactive Visualization of Everyday Memories with Personal Wearable AR Assistant

### v1

@@ Abstract
We introduce Memento, a conversational AR assistant that permanently captures and memorizes user’s verbal queries alongside their spatiotemporal and activity contexts. By storing these “memories,” Memento discovers connections between users’ recurring interests and the contexts that trigger them. Upon detection of similar or identical spatiotemporal activity, Memento proactively recalls user interests and delivers up-to-date responses through AR, seamlessly integrating AR experience into their daily routine. Unlike prior work, each interaction in Memento is not a transient event, but a connected series of interactions with coherent long-term perspective, tailored to the user’s broader multimodal (visual, spatial, temporal, and embodied) context. We conduct preliminary evaluation through user feedbacks with participants of diverse expertise in immersive apps, and explore the value of proactive context-aware AR assistant in everyday settings. We share our findings and challenges in designing a proactive, context-aware AR system.
@@ Introduction
Recent advances in Large Language Models (LLMs) have significantly transformed user interfaces, enabling natural-language conversations that capture the semantics and nuances of user queries, generating rich, context-aware responses. Coupled with AR (Augmented Reality), LLM-powered conversational assistants not only deliver dynamic, real-time answers, but also ground digital contents within the physical environment, creating interactions that empower intuitive engagement with real-world contexts.
@@
Existing systems [chang2024worldscribe, dogan2024augmented, lee2024gazepointar] have demonstrated the potential of combining AR, conversational interface, multi-sensory data, and LLMs to enable immersive, context-aware interactions. These approaches enrich user experiences by overlaying digital content onto everyday environments. However, these systems primarily focus on responding to immediate queries based on current or recent visual inputs, and short-term conversational history. While suitable for generic verbal queries, they fall short of long-term, personalized contexts. Also, the response to a question about an object can vary significantly since the meaning of the object can carry different value across users, and can shift according to spatial (Where), temporal (When), and task-driven (What) contexts. For example, in existing systems, querying about a toy doll would output the cost and the property of the product, regardless of its spatiotemporal context, while it may possess different personal meaning at home–“my kid’s favorite toy”–compare to a store. Thus, a comprehensive understanding of the context behind a user’s query is necessary. Another set of research explores personal “contextual memory banks” that assist users in recalling prior embodied interactions upon queried [li2024omniquery, omniactions, shenencode24]. Yet, there remain opportunities to expand the use-cases beyond user-initiated (reactive) queries, towards proactive, and immersive, situated, visualization of memories.
@@
Moreover, we concentrate on the growing research interest in everyday smart wearable technologies (e.g., HMDs, Glasses, Earbuds, Watches) that lifelog users’ actions and surroundings [arakawa2024prism, fang2025mirai, vinci, shenencode24, yang2025egolife, zulfikar2024memoro]. These works hint at how future personal device form factors can capture patterns and contexts of users’ daily activities, and memorize them for remembrance or insight assistance.
@@
To this end, we introduce Memento, a conversational AR assistant that aims to explore the usability and potential of proactive visualizations combined with contextual memories, on a wearable AR form factor. Memento captures and memorizes a user’s verbal queries (Memento: Towards Proactive Visualization of Everyday Memories with Personal Wearable AR Assistant–“What is the weather like today”, “Tell me the breaking news of the day”), along with their spatiotemporal, activity contexts (Memento: Towards Proactive Visualization of Everyday Memories with Personal Wearable AR Assistant–“House entrance”, “at 7:50AM”, while “Preparing to commute”). Utilizing this personal “contextual memory,” Memento identifies the user’s daily routine, and proactively proposes the visualizations of user’s interests, learned from prior memories. This enables the users to bypass an explicit query amid repeated routines, and serve as an informative visual reminder, as well as a new learning opportunity, within everyday environment.
@@
We pose ‘proactivity’ as one potential interaction pattern of everyday wearable form factors–beyond user-requested unidirectional information flow–given their capability to memorize recurring behaviors and predict users’ needs through multimodal AI-driven assistance [fang2025mirai, liu2024proactive, zulfikar2024memoro]. By designing a prototype proactive context-aware system, Memento, and deploying it in various real-world scenarios, we share preliminary insights of everyday use of a wearable AR. Our main contributions are:
@@
  • We design a proactive AR assistant prototype that situates the visualizations of user’s prior interests (e.g., Weather, Traffic) without explicit requests, serving as a smart visual reminder and live information provider in daily environments.
@@
  • We share the applicability of proactive, context-aware AR assistant system in real-world scenarios via its simplified form and report insights of designing such system.
@@ 1.1 AI and AR Assistance in Routine Daily Life
Recent systems such as MaRginalia [qiu2025marginalia], XR-Objects [dogan2024augmented], and Mirai [fang2025mirai] showed that linking digital content to physical referents and past interactions helps users continue tasks without re-finding information. With emerging wearable XR devices, XR is increasingly positioned as an everyday interface that maintains continuity between physical environments and digital memory [lv2024aria, azuma2019road, barfield2015fundamentals, bressa2022data, tran2025wearable]. Memoro [zulfikar2024memoro] intelligently completes the sentence and assists the user through audio using prior interactions.
The daily use of portable device is highly routine [almourad2024exploring, dengmeasuring, kim2019real, sarker2019, silva2018discovering], yet, current assistants require users to repeat the similar queries, daily, and is reactively triggered. Proactive systems resurface information when contexts recur, enabling continuation of interactions over a restart. Satori [li2025satori] anticipates needs from task-based contexts, and Sensible Agent [lee2025sensible] proactively recognize the needs of user for unobtrusive use of XR in public, and minimized interaction burden.
As XR transitions toward everyday use, proactively resurfacing personal context without repeated requests, aligns with how users revisit routines, spaces, and objects. This motivates treating past interactions as personal “memories” linked to spatial, temporal, and activity cues, and resurfacing them whenever similar routines return, integrating digital recall into daily behavior.
@@ 1.3 Capture and Retrieval of Memories
In situated visualization, the relationship between the user and the interacted physical referents carries significant contextual value. In this work, we define each interaction–capturing spatial, temporal, verbal, visual, and activity context–as a “memory” [barfield2015fundamentals, omar2025, liu2024investigating, satriadi2023proxsituated, satriadi2023context]. Effectively storing and retrieving these memories require robust data management. Traditional techniques such as R-trees and Oct/Quad-trees have been adopted to handle large-scale datasets [koutroumanis2021scalable, qi2020effectively, tian2022survey], and are efficient for geo-tagged data, but struggle with unstructured queries or those requiring semantic context fusion [liu2024suql, caesura]. Recent research leverages vector embeddings from LLMs to represent multimodal data, enabling flexible retrieval through natural language and allowing systems to synthesize semantically coherent responses [asai2023retrieval, chen2022murag, kashmira2024graph, khan2024situational, lewis2020retrieval].
@@ 1.4 User Activity and Context Modeling
Modeling user activity and context is essential for delivering intelligent and timely AR assistance. Prior research leverages multimodal data, including gaze, gesture, speech, and spatial context, to infer user intent [fang2025mirai, vinci, liu2024proactive]. Kim et al. [kim2025explainable] demonstrates that combining user actions, dialogue, and spatiotemporal information enables LLMs to extrapolate intent. Memoro [zulfikar2024memoro] introduces query-less assistance by detecting conversation topics and augmenting user’s remembrance through audio feedback. OmniActions [omniactions] utilizes multi-sensory inputs to anticipate user actions.
Building on these insights, we capture user activities with their personal context to leverage them as a “personal spatial memory”. The memories act as triggers to the information upon semantically aligned context–each spatiotemporal-activity-referent context holds disparate personal memories (Fig. 1).
Figure 1: The view of situated AR in Memento. A physical referent, derives its meaning from its spatiotemporal context (Where and When of the object). An AR content anchored to it serves as an augmentation layer that extends the physically visible information of the referent, within the same context. The referent acts as the bridge between the physical and virtual worlds.
@@ 2 Memento: Proactive Memory-aware Assistant
Figure 2: Memento pipeline overview: At the point of a verbal query, the visual, head-gaze, and the transcribed textual query from the user, is sent to Memento. Upon completion of response generation, the answer is situated onto the referent the query is initiated to, and is permanently stored along with its spatial activity context as a form of a memory. After the initial query, the memory is proactively recalled with the user’s similar contextual spatial activity, and situated onto the referent of user’s prior interest, without an explicit query from the user. Refer to caption Personal Assistant, Refer to caption Spatial Activity Recognizer, Refer to caption Memory Generator, Refer to caption Memory Retriever, Refer to caption Proactive Visualizer, Refer to caption Adjustment Interface
@@ 2.1 Design Considerations
Designing an AR system for everyday use requires principles that support convenience, seamlessness, and robustness across routine contexts. Prior work in ubiquitous and immersive computing highlights the need for systems that move beyond user-initiated interactions toward pervasive assistance [azuma2019road, grubert2016towards]. It shows that proactive support can reduce cognitive effort and improve continuity across repeated behaviors [andolina2018searchbot, liu2024proactive, meurisch2020exploring]. Wearable AR devices further enable embodied, egocentric interaction and situated visualization directly within the user’s viewpoint [lee2023design, yang2025egolife], while practically deploying them in-the-wild requires open-world scene understanding, capable of identifying objects beyond pre-defined data class distribution [Cheng2024YOLOWorld, liu2024grounding, peng2023kosmos]. These insights motivate our design considerations:
Proactive assistance. To reduce repeated explicit queries for routine tasks [andolina2018searchbot, liu2024proactive], we employ semi-proactive resurfacing. An assistance is triggered by recurring user behaviors while allowing dismissal to avoid information overload [liu2023visual, morris2023wearable, zulfikar2024memoro]. A context-aware memory bank stores interaction histories and resurfaces them under matched spatiotemporal and activity cues, with an adjustable interface providing user-agency over repetition cycles [meurisch2020exploring].
@@
Open-world scene analysis. Everyday AR requires recognizing objects users reference, even when they are unseen, during training. To address this, Memento leverages open-vocabulary object detection (YOLO-World [Cheng2024YOLOWorld]) and multimodal LLMs [gemini, gpt4v1] to identify objects from visual and textual descriptions, supporting assistance in unstructured, real-world environments.
While our design lays the groundwork for a full-fledged proactive assistant, in this work, we leverage our prototype system for simplified use-case scenarios. Memento acts as a context-driven reminder that resurfaces relevant information when routines repeat (beyond simple time-based notifications) serving as a step towards long-term proactive, context-aware AR support.
@@ 2.2 Prototype System Design
Implementation Settings. At the time of our prototype development, a raw access to camera sensor (with RGB color) with AR content display, was limited on glasses form factors (e.g., Meta Ray Ban’s), while they are more suitable for daily use of AR. Therefore, we focus on the exploratory use-cases of daily AR assistance than form factor comparisons, and deploy Memento on a Meta Quest3 HMD (with Passthrough AR mode), and collect and discuss user feedbacks for preliminary insight building. Memento is developed with Unity AR Foundation. The HMD is connected to a custom remote Python server, which offloads compute-heavy processes including LLM interactions. Also, the HMD is network-paired with an auxiliary device (smartphone) to provide accurate GPS location of the user. The storing and recalling of memories are crucial internal operations of Memento. We leverage two multimodal LLM models: GPT-4o [gpt4v1] and Gemini 2.0-Flash [gemini], for generating memories, textual descriptions of scene/activity, and referent localization – and CLIP embeddings model [radford2021learning] to vectorize memories, and deterministically compare retrieved memories.
@@ 2.2.1 Personal Conversational Assistant
The Assistant module supports the other five modules of Memento. As shown in Fig. 2, the assistant acts as the starting point to enable a user to interact with a referent in the physical presence. We track the user’s head-gaze and pinch hand gesture to identify the user’s referent of interest. At any point during an AR session, the user can perform the head-gaze-and-pinch to inquire about a topic to the assistant and receive an answer. Upon the detection of a pinch gesture, we capture the user’s view as well as a verbal query, which is transcribed to a text with Whisper model [whisperUnity], and transmitted to the LLM, connected to our server. We also place a visual marker on the 2D view capture, to provide a visual guidance to the referent of user’s query [lee2024gazepointar]. The marker is generated by projecting the 3D position of the referent derived from the gaze, onto the 2D capture. Moreover, to respond to the user’s queries on live information (beyond static knowledge-base of LLMs), we integrate the Google Custom Search API [googlecustomsearch] into the assistant.
@@ 2.2.2 Spatial Activity Analysis
Timely recognizing the user’s current space, and identifying the activity the user is engaged in, is pivotal to Memento. The appropriate inference of these information enables proactive visualizations of context-relevant AR contents. We lifelog the user’s first-person view (via external camera of the HMD), latitude, longitude, and timestamp, in a continuous manner, and infer the user’s current scene and activity. This spatial activity analysis maintains an up-to-date context of the user’s behaviors in the real-world. The interaction contexts includes, the user’s immediate space, activity, time of activity, the referent of the activity, and the geographical location of the user’s location.
A textual description may lack details to fully characterize the user’s spatial property. For example, two locations, both described as “Kitchen” may be in two geographically distinct locations (e.g., kitchens of different households). Thus, we first use GPT-4o to represent the user’s current spatial activity in text, and associate geographical coordinates (latitude, longitude, accuracy: margin-of-error) to the Spatial Activity Profile. This combination of spatial activity contexts (latitude, longitude, visual-text descriptor, activity) ensure Memento to distinguish spaces of the real-world environments, while preserving the fine-grained semantics of each space and activity. The descriptor which is a single vector representation, is generated by prompting GPT-4o to output textual description of the scene given an image capture of the user’s view, and converted to CLIP-encoded embeddings.

### v2

@@ Abstract
We introduce Memento, a conversational AR assistant that permanently captures and memorizes users’ verbal queries alongside their spatiotemporal and activity contexts. By storing these “memories,” Memento discovers connections between users’ recurring interests and the contexts that trigger them. Upon detection of similar or identical spatiotemporal activity, Memento proactively recalls user interests and delivers up-to-date responses through AR, seamlessly integrating the AR experience into their daily routine. Unlike prior work, each interaction in Memento is not a transient event, but part of a connected series of interactions with coherent long-term perspective, tailored to the user’s broader multimodal (visual, spatial, temporal, and embodied) context. We conduct a preliminary evaluation through user feedback from participants of diverse expertise in immersive apps, and explore the value of a proactive context-aware AR assistant in everyday settings. We share our findings and challenges in designing a proactive, context-aware AR system.
@@ Introduction
Recent advances in Large Language Models (LLMs) have significantly transformed user interfaces, enabling natural-language conversations that capture the semantics and nuances of user queries and generating rich, context-aware responses. Coupled with AR (Augmented Reality), LLM-powered conversational assistants not only deliver dynamic, real-time answers, but also ground digital content within the physical environment, creating interactions that empower intuitive engagement with real-world contexts.
@@
Existing systems [chang2024worldscribe, dogan2024augmented, lee2024gazepointar] have demonstrated the potential of combining AR, conversational interface, multi-sensory data, and LLMs to enable immersive, context-aware interactions. These approaches enrich user experiences by overlaying digital content onto everyday environments. However, these systems primarily focus on responding to immediate queries based on current or recent visual inputs, and short-term conversational history. While suitable for generic verbal queries, they fall short of long-term, personalized contexts. Also, the response to a question about an object can vary significantly since the meaning of the object can carry different value across users, and can shift according to spatial (Where), temporal (When), and task-driven (What) contexts. For example, in existing systems, querying about a toy doll would output the cost and the property of the product, regardless of its spatiotemporal context, while it may possess different personal meaning at home–“my kid’s favorite toy”–compared to a store. Thus, a comprehensive understanding of the context behind a user’s query is necessary. Another set of research explores personal “contextual memory banks” that assist users in recalling prior embodied interactions when queried [li2024omniquery, omniactions, shenencode24]. Yet, there remain opportunities to expand the use cases beyond user-initiated (reactive) queries, towards proactive, and immersive, situated visualization of memories.
@@
Moreover, we focus on the growing research interest in everyday smart wearable technologies (e.g., HMDs, glasses, earbuds, watches) that lifelog users’ actions and surroundings [arakawa2024prism, fang2025mirai, vinci, shenencode24, yang2025egolife, zulfikar2024memoro]. These works hint at how future personal device form factors can capture patterns and contexts of users’ daily activities, and memorize them for remembrance or insight assistance.
@@
To this end, we introduce Memento, a conversational AR assistant that aims to explore the usability and potential of proactive visualizations combined with contextual memories on a wearable AR form factor. Memento captures and memorizes a user’s verbal queries (Memento: Towards Proactive Visualization of Everyday Memories with Personal Wearable AR Assistant–“What is the weather like today”, “Tell me the breaking news of the day”), along with their spatiotemporal and activity contexts (Memento: Towards Proactive Visualization of Everyday Memories with Personal Wearable AR Assistant–“House entrance”, “at 7:50AM”, while “Preparing to commute”). Utilizing this personal “contextual memory,” Memento identifies the user’s daily routine and proactively proposes visualizations of the user’s interests learned from prior memories. This enables the users to bypass an explicit query amid repeated routines and serves as an informative visual reminder, as well as a new learning opportunity within everyday environments.
@@
We pose ‘proactivity’ as one potential interaction pattern of everyday wearable form factors–beyond user-requested unidirectional information flow–given their capability to memorize recurring behaviors and predict users’ needs through multimodal AI-driven assistance [fang2025mirai, liu2024proactive, zulfikar2024memoro]. By designing a prototype proactive context-aware system, Memento, and deploying it in various real-world scenarios, we share preliminary insights of everyday use of wearable AR. Our main contributions are:
@@
  • We design a proactive AR assistant prototype that situates visualizations of user’s prior interests (e.g., weather, traffic) without explicit requests, serving as a smart visual reminder and live information provider in daily environments.
@@
  • We share insights on the applicability of proactive, context-aware AR assistant in real-world scenarios via a simplified prototype and report insights into designing such a system.
@@ 1.1 AI and AR Assistance in Routine Daily Life
Recent systems such as MaRginalia [qiu2025marginalia], XR-Objects [dogan2024augmented], and Mirai [fang2025mirai] showed that linking digital content to physical referents and past interactions helps users continue tasks without re-finding information. With emerging wearable XR devices, XR is increasingly positioned as an everyday interface that maintains continuity between physical environments and digital memory [lv2024aria, azuma2019road, barfield2015fundamentals, bressa2022data, tran2025wearable]. Memoro [zulfikar2024memoro] intelligently completes sentences and assists the user through audio using prior interactions.
The daily use of portable devices is highly routine [almourad2024exploring, dengmeasuring, kim2019real, sarker2019, silva2018discovering], yet, current assistants require users to repeat similar queries daily and are reactively triggered. Proactive systems resurface information when contexts recur, enabling continuation of interactions without restarting. Satori [li2025satori] anticipates needs from task-based contexts, and Sensible Agent [lee2025sensible] proactively recognizes the needs of users for unobtrusive use of XR in public and minimizes interaction burden.
As XR transitions toward everyday use, proactively resurfacing personal context without repeated requests aligns with how users revisit routines, spaces, and objects. This motivates treating past interactions as personal “memories” linked to spatial, temporal, and activity cues, and resurfacing them whenever similar routines return, integrating digital recall into daily behavior.
@@ 1.3 Capture and Retrieval of Memories
In situated visualization, the relationship between the user and the physical referents they interact with carries significant contextual value. In this work, we define each interaction–capturing spatial, temporal, verbal, visual, and activity context–as a “memory” [barfield2015fundamentals, omar2025, liu2024investigating, satriadi2023proxsituated, satriadi2023context]. Effectively storing and retrieving these memories requires robust data management. Traditional techniques such as R-trees and Oct/Quad-trees have been adopted to handle large-scale datasets [koutroumanis2021scalable, qi2020effectively, tian2022survey], and are efficient for geo-tagged data, but struggle with unstructured queries or those requiring semantic context fusion [liu2024suql, caesura]. Recent research leverages vector embeddings from LLMs to represent multimodal data, enabling flexible retrieval through natural language and allowing systems to synthesize semantically coherent responses [asai2023retrieval, chen2022murag, kashmira2024graph, khan2024situational, lewis2020retrieval].
@@ 1.4 User Activity and Context Modeling
Modeling user activity and context is essential for delivering intelligent and timely AR assistance. Prior research leverages multimodal data, including gaze, gesture, speech, and spatial context, to infer user intent [fang2025mirai, vinci, liu2024proactive]. Kim et al. [kim2025explainable] demonstrate that combining user actions, dialogue, and spatiotemporal information enables LLMs to extrapolate intent. Memoro [zulfikar2024memoro] introduces query-less assistance by detecting conversation topics and augmenting users’ remembrance through audio feedback. OmniActions [omniactions] utilizes multi-sensory inputs to anticipate user actions.
Building on these insights, we capture user activities with their personal context to leverage them as a “personal spatial memory”. The memories act as triggers for surfacing information when the context is semantically aligned–each spatiotemporal-activity-referent context holds disparate personal memories (Fig. 1).
Figure 1: The view of situated AR in Memento. A physical referent derives its meaning from its spatiotemporal context (Where and When of the object). AR content anchored to it serves as an augmentation layer that extends the physically visible information of the referent, within the same context. The referent acts as the bridge between the physical and virtual worlds.
@@ 2 Memento: Proactive Memory-aware Assistant
Figure 2: Memento pipeline overview: With a verbal query, the visual, head-gaze, and the transcribed textual query from the user are sent to Memento. Upon completion of response generation, the answer is situated onto the referent that the query is initiated to, and is permanently stored along with its spatial activity context as a form of a memory. After the initial query, the memory is proactively recalled when the user’s similar contextual spatial activity, and situated onto the referent of the user’s prior interest, without an explicit query from the user. Refer to caption Personal Assistant, Refer to caption Spatial Activity Recognizer, Refer to caption Memory Generator, Refer to caption Memory Retriever, Refer to caption Proactive Visualizer, Refer to caption Adjustment Interface
@@ 2.1 Design Considerations
Designing an AR system for everyday use requires principles that support convenience, seamlessness, and robustness across routine contexts. Prior work in ubiquitous and immersive computing highlights the need for systems that move beyond user-initiated interactions toward pervasive assistance [azuma2019road, grubert2016towards]. It shows that proactive support can reduce cognitive effort and improve continuity across repeated behaviors [andolina2018searchbot, liu2024proactive, meurisch2020exploring]. Wearable AR devices further enable embodied, egocentric interaction and situated visualization directly within the user’s viewpoint [lee2023design, yang2025egolife], while practically deploying them in-the-wild requires open-world scene understanding capable of identifying objects beyond predefined data class distribution [Cheng2024YOLOWorld, liu2024grounding, peng2023kosmos]. These insights motivate our design considerations:
Proactive assistance. To reduce repeated explicit queries for routine tasks [andolina2018searchbot, liu2024proactive], we employ semi-proactive resurfacing. Assistance is triggered by recurring user behaviors while allowing dismissal to avoid information overload [liu2023visual, morris2023wearable, zulfikar2024memoro]. A context-aware memory bank stores interaction histories and resurfaces them under matched spatiotemporal and activity cues, with an adjustable interface providing users agency over repetition cycles [meurisch2020exploring].
@@
Open-world scene analysis. Everyday AR requires recognizing objects users reference, even when they are unseen during training. To address this, Memento leverages open-vocabulary object detection (YOLO-World [Cheng2024YOLOWorld]) and multimodal LLMs [gemini, gpt4v1] to identify objects from visual and textual descriptions, supporting assistance in unstructured, real-world environments.
While our design lays the groundwork for a full-fledged proactive assistant, in this work, we leverage our prototype system for simplified use case scenarios. Memento acts as a context-driven reminder that resurfaces relevant information when routines repeat (beyond simple time-based notifications) serving as a step towards long-term proactive, context-aware AR support.
@@ 2.2 Prototype System Design
Implementation Settings. At the time of our prototype development, raw access to a camera sensor (with RGB color) with AR content display, was limited on glasses form factors (e.g., Meta Ray Ban’s), while these form factors are more suitable for daily use of AR. Therefore, we focus on the exploratory use-cases of daily AR assistance rather than form factor comparisons, and deploy Memento on a Meta Quest3 HMD (with Passthrough AR mode), and collect and discuss user feedback for preliminary insight building. Memento is developed with Unity AR Foundation. The HMD is connected to a custom remote Python server, which offloads compute-heavy processes including LLM interactions. In addition, the HMD is network-paired with an auxiliary device (smartphone) to provide accurate GPS location of the user. The storing and recalling of memories are crucial internal operations of Memento. We leverage two multimodal LLM models: GPT-4o [gpt4v1] and Gemini 2.0-Flash [gemini], for generating memories, textual descriptions of scenes/activities, and referent localization – and CLIP embedding model [radford2021learning] to vectorize memories, and deterministically compare retrieved memories.
@@ 2.2.1 Personal Conversational Assistant
The Assistant module supports the other five modules of Memento. As shown in Fig. 2, the assistant acts as the starting point that enables a user to interact with a referent in the physical environment. We track the user’s head-gaze and pinch gesture to identify the user’s referent of interest. At any point during an AR session, the user can perform the head-gaze-and-pinch to inquire about a topic from the assistant and receive an answer. Upon the detection of a pinch gesture, we capture the user’s view as well as a verbal query, which is transcribed into text with the Whisper model [whisperUnity], and transmitted to the LLM, connected to our server. We also place a visual marker on the 2D view capture to provide visual guidance to the referent of the user’s query [lee2024gazepointar]. The marker is generated by projecting the 3D position of the referent derived from the gaze, onto the 2D capture. Moreover, to respond to the user’s queries with live information (beyond the static knowledge-base of LLMs), we integrate the Google Custom Search API [googlecustomsearch] into the assistant.
@@ 2.2.2 Spatial Activity Analysis
Timely recognizing the user’s current space, and identifying the activity the user is engaged in, is pivotal to Memento. The appropriate inference of this information enables proactive visualizations of context-relevant AR content. We lifelog the user’s first-person view (via the external camera of the HMD), latitude, longitude, and timestamp, in a continuous manner, and infer the user’s current scene and activity. This spatial activity analysis maintains an up-to-date context of the user’s behaviors in the real-world. The interaction contexts include, the user’s immediate space, activity, time of activity, the referent of the activity, and the geographical location of the user.
A textual description may lack sufficient details to fully characterize the user’s spatial context. For example, two locations, both described as “Kitchen” may be in two geographically distinct locations (e.g., kitchens of different households). Thus, we first use GPT-4o to represent the user’s current spatial activity in text, and associate geographical coordinates (latitude, longitude, accuracy: margin-of-error) with the Spatial Activity Profile. This combination of spatial activity contexts (latitude, longitude, visual-text descriptor, activity) enables Memento to distinguish spaces of the real-world spaces, while preserving the fine-grained semantics of each space and activity. The descriptor which is a single vector representation, is generated by prompting GPT-4o to output textual description of the scene given an image capture of the user’s view, and converted to CLIP-encoded embeddings.